an open pipe resonates with frequency f1 and 2nd harmonic. Now one end is closed and the frequency is slowly increased till it resonates with frequency f2 and nth harmonic, then n=? and find the relation between f1 and and f2.


P.S.- after solving my answer comes to n=3 and f2=3/5f1 but the books i have referred to show the answer to be n=5 and f2=5/4f1

Grade:12

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

f1=1/l root(B/p)

f2=n/4l root(B/p)

f2=n/4f1

for the 1st resoance n=5

f2=5/4 f1

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi